OpenStudy (anonymous):

Find the equation of the parabola that has a vertex of (4,1) and passes through the point (3,0)

OpenStudy (amistre64):

since you know the vertex, the vertex form would seem a good start

OpenStudy (amistre64):

y = a(x-vx)+vy ; where vertex = (vx,vy)

OpenStudy (amistre64):

forgot my ^2

OpenStudy (amistre64):

lets fill in the specifics now: y = a(x-4)^2+1 now, when x=3, and y=0; what do we get to work with? 0 = a(3-4)^2 +1 0 = a(-1)^2 +1 0 = a +1 -1 = a

OpenStudy (amistre64):

so write it up with the proper a value and it should be good. you might have to expand it to match a choice